How many pounds of peanuts costing $3.00 a pound should be mixed with 4 pounds of cashews costing $4.50 a pound to obtain a mixture costing $3.50 a pound?
Answer by stanbon(75887)

How many pounds of peanuts costing $3.00 a pound should be mixed with 4 pounds of cashews costing $4.50 a pound to obtain a mixture costing $3.50 a pound?
----------------
Equation::
value + value = value
3x + 4*4.5 = 3.5(x+4.5)
3x + 4*4.5 = 3.5x + 3.5*4.5
-----
0.5x = 0.5*4.5
---
x = 4.5 lbs (amt. of $3.00 peanuts needed)
-------------
Cheers,
Stan H.
-------------